Hall & Company issues $10,000,000 in 4% bonds with detachable warrants of 12 warrants for each $1,000 bond.  The bonds are issued at 99.   Record the issuance under the following assumptions:   Assume that the fair value of the warrants are $6.00 or $720,000 and the bonds would have been issued at 96.
Answer #1

a) Market Value of bonds= 10000*96= $960,000 ($10,000,000/$1,000)

b) Cost of warrants: 10000 warrants * $6 (fair value) *12 (detachable warrants)=$720,000

Therefore, total of a and b = $1,680,000.

% of a to total: $960,000/$1,680,000 * total cash proceeds: ~5657140 (Approx)

The bonds are issued at a discount of 1% therefore cash proceeds will be $10,000,000*0.99=$9,900,000. This amount will become the debit entry.

Cash Acc $9,900,000